# Descent-Guided Policy Gradient for Scalable Cooperative Multi-Agent Learning

### v1

@@ 6.3 Controlled Comparison
Figure 3 compares training dynamics at $N\in\{2,5,10\}$. At N=2N{=}2, DG-PG converges to near-Best-Fit performance within 5 episodes and remains stable, while MAPPO and IPPO briefly approach similar levels but subsequently deteriorate. At N=5N{=}5, the advantage becomes more pronounced: DG-PG achieves a mean reward of 14.2-14.2 (close to Best-Fit’s 12.6-12.6), while MAPPO (34.1-34.1) and IPPO (39.8-39.8) plateau at $2\text{--}3\times$ worse performance. At N=10N{=}10, IPPO deteriorates sharply (49.3-49.3), confirming that independent learning fails as coordination demands grow. In contrast, DG-PG reaches 11.7-11.7, surpassing the Best-Fit heuristic through temporal strategies unavailable to the single-step greedy policy. These results confirm that gradient guidance is the decisive factor; the natural next question is whether this advantage persists at larger scales.
@@ 6.4 Scalability
Figure 5 overlays DG-PG’s training curves across all scales. A striking finding emerges: DG-PG converges to within 10% of its best observed reward in $\leq$10 episodes at every scale—from N=5N{=}5 to N=200N{=}200. The near-perfect alignment of the six curves provides direct empirical confirmation of the agent-independent sample complexity predicted by Theorem 4.3: the number of episodes required for convergence does not grow with NN.
@@
DG-PG’s rapid convergence produces policies that generalize well beyond training. Performance remains stable as $\alpha$ decays from 0.9 to 0.2, indicating that agents internalize effective coordination patterns early and do not depend on the guidance signal to maintain them. Unlike residual policy methods, which require the base controller to remain online at deployment, DG-PG uses the analytical reference only during training; at test time the policy runs independently. On 40 held-out test scenarios—where the guidance signal is entirely absent—the resulting decentralized policy matches or exceeds Best-Fit, a centralized heuristic with access to full global state, across the majority of tested scales (Table 1).
@@ 7 Conclusion
In this work, we identified cross-agent noise—the interference that each agent’s learning signal receives from the stochastic exploration of all other agents—as the central bottleneck in scaling cooperative MARL. DG-PG addresses this by deriving a per-agent guidance signal from a domain-specific analytical reference and using it to reshape the policy gradient. This reduces gradient variance from $\Theta(N)$ to $\mathcal{O}(1)$ while preserving the Nash equilibria of the original game. Experiments on a heterogeneous cloud scheduling benchmark confirm that DG-PG scales cooperative learning to 200 agents, a regime where MAPPO and IPPO cannot learn.
More broadly, DG-PG requires only that the analytical reference provide informative gradient directions, not that it be accurate enough to serve as a deployable policy. This makes the approach applicable to any cooperative domain where approximate analytical models are available, such as traffic networks and supply chains.
Limitations and future work. DG-PG is applicable to cooperative domains where an analytical model with informative gradients is available; problems without such structure need alternative approaches. The guidance weight $\alpha$ follows a fixed schedule independent of $N$; scale-adaptive scheduling that increases guidance at larger $N$ is a promising direction to close the remaining gap at N=200N{=}200.
@@ Appendix A Comparison with Existing Methods
Table 2 provides a systematic comparison of DG-PG with representative methods from major paradigms in cooperative MARL. We evaluate methods across key dimensions: gradient variance scaling, sample complexity, support for heterogeneous policies, parallel updates, architectural overhead, and domain knowledge requirements. The comparison highlights DG-PG’s unique position as the only method that simultaneously achieves $\mathcal{O}(1)$ gradient variance, supports heterogeneous agent policies with parallel updates, and requires no additional learned components beyond a standard centralized critic.
@@
QPLEX (Wang et al., 2021a) Attention N/A (Value) Attn mixer, $\mathcal{O}(N^{2})$ Dueling struct.
@@
Diff. Rewards (Wolpert and Tumer, 2001) Counterfactual $\mathcal{O}(1)^{*}$ $\mathcal{O}(1/\epsilon)^{*}$ None Exact simulator
@@
HAPPO reduces per-agent variance via sequential updates with trust-region constraints, but sacrifices parallelism: wall-clock time per iteration scales as $\mathcal{O}(N)$, yielding effective sample complexity $\mathcal{O}(N/\epsilon)$ in wall-clock time.
@@
Difference Rewards achieve $\mathcal{O}(1)$ variance in principle but require an exact reward function or high-fidelity simulator to compute $r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})$ for each agent, which is often intractable at scale.
§Mean-Field results hold under the mean-field approximation (agent exchangeability). Convergence is to the mean-field equilibrium, which may differ from the Nash equilibrium of the original $N$-agent game.
@@
DG-PG satisfies all five simultaneously (Theorems 4.24.3), with only a minimal modification to the advantage computation in standard MAPPO.
2. Leveraging analytical models without sacrificing optimality. All other methods in Table 2 are purely model-free: they either impose no assumptions at all (MAPPO, IPPO, HAPPO, COMA) or encode structural constraints into the architecture (QMIX, VDN, Mean-Field). DG-PG is unique in leveraging an external analytical model to reduce variance while remaining model-free in its convergence guarantees. This distinguishes it from other knowledge-informed learning paradigms: Physics-Informed Neural Networks (Raissi et al., 2019) embed governing equations directly into the loss, requiring exact PDEs; Residual RL (Johannink et al., 2019) uses a hand-designed controller as a permanent base policy, inheriting its biases; Guided Policy Search (Levine and Koltun, 2013) relies on trajectory optimization with known dynamics. In contrast, DG-PG treats the analytical model purely as a control variate—its gradient contribution is subtracted out in expectation, so model inaccuracies affect only variance, not the stationary points of the objective (Theorem 4.1). The adjustable $\alpha$ further allows practitioners to tune the degree of reliance based on model fidelity.
@@ Appendix B Theoretical Assumptions and Dependencies
Thm. 4.1 (Nash Invariance) Exogeneity ensures reference does not shift; Alignment ensures guidance vanishes at optimality.
Thm. 4.2 (Variance Bound) Exogeneity makes reference deterministic; Alignment ensures positive correlation $\rho>0$.
Thm. 4.3 (Sample Complexity) Smoothness enables descent lemma; PL condition yields linear convergence rate; $\mathcal{O}(1)$ variance from Thm. 4.2.
@@
  • Assumption 3.1 (Exogeneity) is a design choice rather than a restrictive condition: it is satisfied by construction when the reference is computed from exogenous system parameters (e.g., server capacities, total workload) and treated as a fixed target during gradient computation via the stop-gradient operator $\bot(\cdot)$.
@@ Appendix C Case Study: Verification of Structural Assumptions
In each case, the key requirement is that the prior defines a descent direction aligned with the cooperative objective—a property we formally verify below for the cloud scheduling domain using fluid-limit analysis from queueing theory (Kurtz, 1981).
@@ C.1 System Model and Objective Approximation
Surrogate Objective via Queueing Theory. To construct the reference state $\tilde{\boldsymbol{X}}$, we need an analytically tractable objective whose minimizer serves as the guidance target. The cooperative RL objective $J$ itself is intractable—it aggregates queueing delay, energy efficiency, and fairness through complex, non-separable interactions. However, queueing theory reveals that improving each of these criteria reduces to a common structural condition:
@@
Equivalently, the optimal state satisfies the utilization balancing condition: the utilization rate $x^{i,k}/\mu^{i,k}$ is equalized across all servers for each resource type $k$:
@@
Generalization to Other Domains. The verification procedure demonstrated above—constructing a convex surrogate objective, deriving its optimal solution as the reference, and proving descent alignment via first-order optimality conditions—serves as a template for applying DG-PG to other systems. As outlined in the main text, domains such as communication networks (proportional fairness), power grids (economic dispatch), and traffic systems (Wardrop equilibrium) all admit analytical priors with similar structural properties. In each case, the key steps are: (1) identify a domain-specific performance metric with known convexity properties; (2) derive the reference state as its constrained optimum; (3) verify that the reference defines a descent direction using tools from convex analysis or variational inequalities. This case study establishes the feasibility and rigor of such verification, providing a blueprint for extending DG-PG beyond cloud scheduling.
Robustness to Approximate Satisfaction. In practice, the structural assumptions may not hold exactly. For instance, the fluid-limit approximation underlying the reference state assumes Poisson arrivals and homogeneous service times, which are often violated in real systems due to heavy-tailed workloads and temporal fluctuations. However, DG-PG remains effective as long as the assumptions hold approximately—that is, the reference captures the dominant performance drivers even if the analytical model is idealized. The adjustable guidance weight $\alpha$ provides a natural mechanism for handling model inaccuracies: when the prior is less accurate, reducing $\alpha$ diminishes its influence while preserving the variance reduction benefits. Our experiments (Section 6.2) demonstrate that DG-PG is robust across a range of $\alpha$ values, confirming that directional correctness, rather than exact model fidelity, is sufficient for effective guidance.
@@ Appendix D Derivation of Local Influence Vector
In this appendix, we derive the local influence vector $\boldsymbol{z}_{t}^{i}=\frac{\partial\boldsymbol{x}_{t}}{\partial a^{i}_{t}}$ for the cloud resource scheduling example. This derivation generalizes to other domains with similar additive structure.
We use the notation established in Appendix C: $N$ servers, $K$ resource dimensions, global utilization vector $\boldsymbol{x}_{t}\in\mathbb{R}^{NK}$, and per-server utilization $\boldsymbol{x}_{t}^{j}\in\mathbb{R}^{K}$. Each agent $i$ holds a task with resource requirement vector $\boldsymbol{w}^{i}=(w^{i,1},\dots,w^{i,K})^{\top}\in\mathbb{R}^{K}$, where $w^{i,k}$ is the demand on resource dimension $k$ (e.g., CPU cores, memory). Agent $i$ selects a target server $a_{t}^{i}=j$, updating server $j$’s utilization as:
$\boldsymbol{x}_{t+1}^{j}=\boldsymbol{x}_{t}^{j}+\boldsymbol{w}^{i}\cdot\mathbb{I}(a_{t}^{i}=j).$
Since only server $j$’s utilization changes, differentiating with respect to $a_{t}^{i}$ yields a sparse influence vector:
$\boldsymbol{z}_{t}^{i}=(\mathbf{0},\dots,\underbrace{\boldsymbol{w}^{i}}_{\text{server }j},\dots,\mathbf{0})^{\top}.$
@@
Projected coefficient. In the DG-PG update (Equations 910), each agent’s guidance signal is scaled by a coefficient $c_{t}^{i}=\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\boldsymbol{z}_{t}^{i}\rangle$ that measures how much agent $i$’s action contributes to the system’s deviation from the reference. Substituting $\boldsymbol{z}_{t}^{i}$ yields:
$c_{t}^{i}=\sum_{k=1}^{K}w^{i,k}\cdot(x_{t}^{j,k}-\tilde{x}_{t}^{j,k}).$
Each term measures how overloaded server $j$ is on resource $k$, weighted by the task’s own demand $w^{i,k}$—so resource types that the task consumes most heavily dominate the signal. Importantly, $c_{t}^{i}$ is a gradient signal that provides a direction for policy improvement, not a hard constraint that forces convergence to $\tilde{\boldsymbol{X}}$—the actual system trajectory is determined by the learned policy interacting with the environment.
@@ D.1 General Applicability
  • Traffic routing: $\boldsymbol{x}_{t}$ as link flows, actions as path selections; $\boldsymbol{z}_{t}^{i}$ indicates which links are used.
  • Inventory management: $\boldsymbol{x}_{t}$ as stock levels, actions as replenishment orders; $\boldsymbol{z}_{t}^{i}$ adds to specific stock dimensions.
  • Communication networks: $\boldsymbol{x}_{t}$ as channel interference levels, actions as power allocations.
@@
In this appendix, we provide the detailed proof for Theorem 4.1, establishing that our DG-PG framework preserves the stationary points of the original cooperative game.
Theorem 4.1 (Restated). Let $\boldsymbol{\theta}^{*}$ be a stationary point of the original cooperative objective, satisfying $\nabla_{\boldsymbol{\theta}}J(\boldsymbol{\theta}^{*})=\mathbf{0}$. Under Assumptions 3.1 and 3.2, $\boldsymbol{\theta}^{*}$ remains a stationary point of the regularized objective $J_{\alpha}$ for any $\alpha\in(0,1)$.
To prove this theorem, we first establish a key lemma stating that the guidance gradient vanishes at any stationary point of the original objective.

### v2

@@ 6.3 Controlled Comparison
Figure 3 compares training dynamics at $N\in\{2,5,10\}$. At N=2N{=}2, DG-PG converges to near-Best-Fit performance within 5 episodes and remains stable, while MAPPO and IPPO briefly approach similar levels but subsequently deteriorate. At N=5N{=}5, the advantage becomes more pronounced, with DG-PG achieving a mean reward of 14.2-14.2 (close to Best-Fit’s 12.6-12.6), while MAPPO (34.1-34.1) and IPPO (39.8-39.8) plateau at $2\text{--}3\times$ worse performance. At N=10N{=}10, IPPO deteriorates sharply (49.3-49.3), confirming that independent learning fails as coordination demands grow. In contrast, DG-PG reaches 11.7-11.7, surpassing the Best-Fit heuristic through temporal strategies unavailable to the single-step greedy policy. These results confirm that gradient guidance is the decisive factor; the natural next question is whether this advantage persists at larger scales.
@@ 6.4 Scalability
Figure 5 overlays DG-PG’s training curves across all scales. Notably, DG-PG converges to within 10% of its best observed reward in $\leq$10 episodes at every scale, from N=5N{=}5 to N=200N{=}200. The near-perfect alignment of the six curves provides direct empirical confirmation of the agent-independent sample complexity predicted by Theorem 4.3: the number of episodes required for convergence does not grow with NN.
@@
DG-PG’s rapid convergence produces policies that generalize well beyond training. Performance remains stable as $\alpha$ decays from 0.9 to 0.2, indicating that agents internalize effective coordination patterns early and do not depend on the guidance signal to maintain them. Unlike residual policy methods, which require the base controller to remain online at deployment, DG-PG uses the analytical reference only during training. At test time, the policy runs independently without guidance. On 40 held-out test scenarios, where the guidance signal is entirely absent, the resulting decentralized policy matches or exceeds Best-Fit, a centralized heuristic with access to full global state, across the majority of tested scales (Table 1).
@@ 7 Conclusion
In this work, we identified cross-agent noise as the central bottleneck in scaling cooperative MARL. DG-PG addresses this by augmenting each agent’s policy gradient with an analytically computed per-agent guidance term derived from a domain-specific reference model. This reduces gradient variance from $\Theta(N)$ to $\mathcal{O}(1)$ while preserving the Nash equilibria of the original game. Experiments on a heterogeneous cloud scheduling benchmark confirm that DG-PG maintains consistent performance from 5 to 200 agents, while standard multi-agent policy gradient methods degrade rapidly with NN.
DG-PG requires only that the analytical model indicate a beneficial direction for the system state, not that the reference be exactly achievable or directly translatable into agent actions. This makes DG-PG applicable to any cooperative domain where approximate analytical models are available, such as traffic networks and supply chains.
Limitations and future work. DG-PG is applicable to cooperative domains where an analytical model with informative gradients is available, and problems without such structure need alternative approaches. The guidance weight $\alpha$ follows a fixed schedule independent of $N$. Scale-adaptive scheduling that increases guidance at larger $N$ is a promising direction to close the remaining gap at N=200N{=}200.
@@ Appendix A Comparison with Existing Methods
Table 2 provides a systematic comparison of DG-PG with representative methods from major paradigms in cooperative MARL. We evaluate methods across key dimensions, including gradient variance scaling, sample complexity, support for heterogeneous policies, parallel updates, architectural overhead, and domain knowledge requirements. The comparison highlights DG-PG’s unique position as the only method that simultaneously achieves $\mathcal{O}(1)$ gradient variance, supports heterogeneous agent policies with parallel updates, and requires no additional learned components beyond a standard centralized critic.
@@
QPLEX (Wang et al., 2021) Attention N/A (Value) Attn mixer, $\mathcal{O}(N^{2})$ Dueling struct.
@@
Diff. Rewards (Wolpert and Tumer, 2001) Counterfactual Problem-dep. Problem-dep. None Exact simulator
@@
HAPPO reduces per-agent variance via sequential updates with trust-region constraints, but sacrifices parallelism. Consequently, wall-clock time per iteration scales as $\mathcal{O}(N)$, yielding effective sample complexity $\mathcal{O}(N/\epsilon)$ in wall-clock time.
@@
Difference Rewards reduce variance by subtracting a counterfactual baseline $r(s,\boldsymbol{a})-r(s,\boldsymbol{a}^{-i})$, but the degree of reduction depends on the reward structure (separability). For non-separable rewards, cross-agent interaction terms persist. Computing the counterfactual also requires an exact simulator, which is often intractable at scale.
§Mean-Field methods assume homogeneous agents (identical and exchangeable). Results hold only under this approximation, and convergence is to the mean-field equilibrium, which may differ from the Nash equilibrium of the original $N$-agent game.
@@
DG-PG satisfies all five simultaneously (Theorems 4.24.3), with only a minimal modification to the advantage computation in standard policy gradient algorithm such as MAPPO.
2. Leveraging analytical models without sacrificing optimality. All other methods in Table 2 are purely model-free, either imposing no assumptions at all (MAPPO, IPPO, HAPPO, COMA) or encoding structural constraints into the architecture (QMIX, VDN, Mean-Field). DG-PG is unique in leveraging an external analytical model to reduce variance while remaining model-free in its convergence guarantees.
DG-PG’s gradient-level use of analytical models also distinguishes it from other knowledge-informed learning paradigms: Physics-Informed Neural Networks (Raissi et al., 2019) embed governing equations directly into the loss, requiring exact PDEs; Residual RL (Johannink et al., 2019) uses a hand-designed controller as a permanent base policy, inheriting its biases; Guided Policy Search (Levine and Koltun, 2013) relies on trajectory optimization with known dynamics. These methods all incorporate domain knowledge into the objective or policy. In contrast, DG-PG derives the gradient of a guidance term from the analytical model to steer policy updates, leaving the original objective $J$ unchanged. As a result, model inaccuracies affect only variance, not the Nash equilibria of the original game (Theorem 4.1). The adjustable $\alpha$ further allows practitioners to tune the degree of reliance based on model fidelity.
@@ Appendix B Theoretical Assumptions and Dependencies
Thm. 4.1 (Nash Invariance) Exogeneity ensures reference does not shift; Alignment ensures guidance gradient vanishes at Nash equilibria.
Thm. 4.2 (Variance Bound) Exogeneity makes reference deterministic; Alignment ensures negative correlation $\rho<0$.
Thm. 4.3 (Sample Complexity) Smoothness enables the per-iteration bound; PL condition yields linear convergence rate; $\mathcal{O}(1)$ variance from Thm. 4.2.
@@
  • Assumption 3.1 (Exogeneity) is a design choice rather than a restrictive condition. It is satisfied by construction when the reference is computed from exogenous system parameters (e.g., server capacities, total workload) and treated as a fixed target during gradient computation via the stop-gradient operator $\bot(\cdot)$.
@@ Appendix C Case Study: Verification of Structural Assumptions
In each case, the key requirement is that the prior defines a descent direction aligned with the cooperative objective, a property we formally verify below for the cloud scheduling domain using fluid-limit analysis from queueing theory (Kurtz, 1981).
@@ C.1 System Model and Objective Approximation
Surrogate Objective via Queueing Theory. To construct the reference state $\tilde{\boldsymbol{X}}$, we need an analytically tractable objective whose minimizer serves as the guidance target. The cooperative MARL objective $J$ itself is intractable because it aggregates queueing delay, energy efficiency, and fairness through complex, non-separable interactions. However, queueing theory reveals that improving each of these criteria reduces to a common structural condition:
@@
Equivalently, the optimal state satisfies the utilization balancing condition, i.e., the utilization rate $x^{i,k}/\mu^{i,k}$ is equalized across all servers for each resource type $k$:
@@
Generalization to Other Domains. The verification procedure demonstrated above serves as a template for applying DG-PG to other systems. As outlined in the main text, domains such as communication networks (proportional fairness), power grids (economic dispatch), and traffic systems (Wardrop equilibrium) all admit analytical priors with similar structural properties. In each case, the key steps are: (1) identify a domain-specific performance metric with known convexity properties; (2) derive the reference state as its constrained optimum; (3) verify that the reference defines a descent direction using tools from convex analysis or variational inequalities. This case study establishes the feasibility and rigor of such verification, providing a blueprint for extending DG-PG beyond cloud scheduling.
Robustness to Approximate Satisfaction. In practice, the structural assumptions may not hold exactly. For instance, the fluid-limit approximation underlying the reference state assumes Poisson arrivals and homogeneous service times, which are often violated in real systems due to heavy-tailed workloads and temporal fluctuations. However, DG-PG remains effective as long as the assumptions hold approximately, that is, the reference captures the dominant performance drivers even if the analytical model is idealized. The adjustable guidance weight $\alpha$ provides a natural mechanism for handling model inaccuracies. When the prior is less accurate, reducing $\alpha$ diminishes its influence while preserving the variance reduction benefits. Our experiments (Section 6.2) demonstrate that DG-PG is robust across a range of $\alpha$ values, confirming that directional correctness, rather than exact model fidelity, is sufficient for effective guidance.
@@ Appendix D Derivation of Local Influence Vector
In this appendix, we derive the local influence vector $\boldsymbol{z}_{t}^{i}=\frac{\partial\boldsymbol{x}_{t}}{\partial a^{i}_{t}}$ for the cloud resource scheduling example. Here $\boldsymbol{z}_{t}^{i}$ captures the marginal effect of agent $i$’s action on the system state, i.e., how choosing a particular server changes the global utilization vector. This derivation generalizes to other domains with similar additive structure.
We use the notation established in Appendix C: $N$ servers, $K$ resource dimensions, global utilization vector $\boldsymbol{x}_{t}\in\mathbb{R}^{NK}$, and per-server utilization $\boldsymbol{x}_{t}^{j}\in\mathbb{R}^{K}$. Each agent $i$ holds a task with resource requirement vector $\boldsymbol{w}^{i}=(w^{i,1},\dots,w^{i,K})^{\top}\in\mathbb{R}^{K}$, where $w^{i,k}$ is the demand on resource dimension $k$ (e.g., CPU cores, memory). Each agent $i$ selects a target server $a_{t}^{i}\in\{1,\dots,N\}$, and the next-step utilization on server $j$ is determined by the joint actions of all agents:
$\boldsymbol{x}_{t+1}^{j}=\boldsymbol{x}_{t}^{j}+\sum_{l=1}^{N}\boldsymbol{w}^{l}\cdot\mathbb{I}(a_{t}^{l}=j).$
To compute the marginal effect of agent $i$’s action, we differentiate with respect to $a_{t}^{i}$. Since the contributions of all other agents $l\neq i$ are held fixed, their terms vanish, leaving only agent $i$’s contribution:
$\boldsymbol{z}_{t}^{i}\triangleq\frac{\partial\boldsymbol{x}_{t+1}}{\partial a_{t}^{i}}=(\mathbf{0},\dots,\underbrace{\boldsymbol{w}^{i}}_{\text{server }j},\dots,\mathbf{0})^{\top}.$
@@
Projected coefficient. In the DG-PG update (Equations 810), each agent’s guidance signal is scaled by the inner product $\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\boldsymbol{z}_{t}^{i}\rangle$, which measures how much agent $i$’s action contributes to the change in deviation from the reference. Substituting the influence vector $\boldsymbol{z}_{t}^{i}$ yields:
$\langle\boldsymbol{x}_{t}-\tilde{\boldsymbol{x}}_{t},\boldsymbol{z}_{t}^{i}\rangle=\sum_{k=1}^{K}w^{i,k}\cdot(x_{t}^{j,k}-\tilde{x}_{t}^{j,k}).$
Each term measures how overloaded server $j$ is on resource $k$, weighted by the task’s own demand $w^{i,k}$, so resource types that the task consumes most heavily dominate the signal. Importantly, this coefficient provides a direction for policy improvement, not a hard constraint that forces convergence to $\tilde{\boldsymbol{X}}$. The actual system trajectory is determined by the learned policy interacting with the environment.
@@ D.1 General Applicability
  • Traffic routing: $\boldsymbol{x}_{t}$ represents link flows and actions are path selections, so $\boldsymbol{z}_{t}^{i}$ indicates which links are used.
  • Inventory management: $\boldsymbol{x}_{t}$ represents stock levels and actions are replenishment orders, where $\boldsymbol{z}_{t}^{i}$ adds to specific stock dimensions.
  • Communication networks: $\boldsymbol{x}_{t}$ represents channel interference levels and actions are power allocations, and $\boldsymbol{z}_{t}^{i}$ captures each agent’s marginal contribution to interference.
@@
In this appendix, we provide the detailed proof for Theorem 4.1, establishing that our DG-PG framework preserves the Nash equilibria of the original cooperative game.
Theorem 4.1 (Nash Invariance). Let $\boldsymbol{\theta}^{*}$ be a Nash equilibrium of the original cooperative game, i.e., $\nabla_{\theta^{i}}J(\boldsymbol{\theta}^{*})=\mathbf{0}$ for all $i\in\mathcal{N}$. Under Assumptions 3.13.2, the DG-PG gradient also vanishes at $\boldsymbol{\theta}^{*}$, i.e., $\nabla_{\theta^{i}}J_{\alpha}(\boldsymbol{\theta}^{*})=\mathbf{0}$ for any $\alpha\in(0,1)$ and all $i\in\mathcal{N}$.
To prove this theorem, we first establish a key lemma stating that the gradient of the guidance term vanishes at any Nash equilibrium of the original game.